# The Confidence Paradox: Can LLM Know When It’s Wrong?

### v1

@@ 5.2 Cross-Domain Generalization Testing
To assess the robustness and generalization ability of HonestVQA framework, we conduct a series of cross-domain testing experiments. These experiments evaluate whether the hallucination detection model trained on one dataset can effectively identify hallucinations in DocVQA outputs on different datasets. Such generalization is crucial in real-world scenarios where the distribution of questions and visual-textual content varies significantly across domains such as scanned documents, infographics, and structurally diverse textual scenes. We train HonestVQA on the source dataset and evaluate it on a different target dataset without any further fine-tuning, measuring hallucination detection performance using Accuracy and F1-score as the primary metrics. From Table 3, we observe that HonestVQA generalizes robustly across domain shifts. Notably, when trained on InfographicsVQA and evaluated on SpDocVQA, the model achieves a high F1-score of 76.1%, outperforming the reverse setting (SpDocVQA $\rightarrow$ InfographicsVQA), which yields 71.8%. This suggests that the high-density, information-rich visual patterns in InfographicsVQA provide transferable inductive biases that enhance hallucination detection in more structured domains like documents. Similarly, the SROIE $\rightarrow$ SpDocVQA setup results in an F1 of 72.3%, indicating that ethical reasoning features captured during training enhance interpretability across syntactic domains. However, the model exhibits relatively lower performance when transferring from SpDocVQA to SROIE (F1 = 69.8%), highlighting the challenges posed by ethical hallucination detection under unfamiliar structural constraints. Nonetheless, the use of uncertainty-aware calibration via confidence-alignment and contrastive ethical loss contributes to soft regularization of decision boundaries, allowing for improved generalization even in low-overlap semantic settings. Interestingly, models trained on SROIE also generalize well to visually and semantically distinct domains such as InfographicsVQA, achieving 67.2% F1. This supports our hypothesis that the inclusion of contrastive ethical supervision enforces more generalizable representations. Furthermore, the relative drop in performance in domain transfer settings (typically within 4–6% of in-domain results) underscores the importance of calibration-aware models in mitigating performance degradation due to domain shift.
@@
Table 5: Sensitivity analysis of alignment and contrastive hyperparameters on SpDocVQA, InfographicsVQA, and SROIE. Note: Accuracy (Acc) and H-Score (H) are used as abbreviated. Bold rows indicate optimal settings.
@@ 5.3 Ablation Study
To thoroughly evaluate the contribution of individual modules in HonestVQA, we conduct an ablation study across three datasets by systematically disabling the confidence-accuracy alignment loss and the contrastive ethical enforcement loss. Table 4 shows that the removal of either module consistently degrades performance across all datasets in terms of accuracy and H-Score, confirming their complementary roles. For instance, on SpDocVQA, the full model achieves 75.9% accuracy and an H-Score of 0.113. Removing the alignment loss drops accuracy to 72.1% and worsens H-Score to 0.172. On InfographicsVQA, the full model yields 68.3% accuracy and an H-Score of 0.134, whereas removing contrastive enforcement lowers accuracy to 65.2% and degrades H-Score to 0.189. Similar trends are observed on SROIE, where the full model achieves 80.2% accuracy and 0.096 H-Score, significantly outperforming the ablated variants. We further conduct a hyperparameter sensitivity analysis on alignment weight $\alpha$, contrastive margin $m$, and loss weights $\lambda_{1}$ and $\lambda_{2}$ across three datasets and summarize the trends in Table 5. The model maintains high performance for $\alpha$ between 0.5 and 1.5, margin $m$ from 0.3 to 0.7, and loss weights $\lambda_{1}=0.1$, $\lambda_{2}=0.05$. Deviations outside these ranges result in decreased accuracy or calibration degradation.
@@ 6.2 Computational Analysis
Practical deployment of DocVQA systems, particularly those incorporating hallucination detection like HonestVQA, requires a thorough understanding of computational costs and resource demands. To assess the efficiency of HonestVQA relative to baseline models, we conduct a detailed latency and resource usage analysis focusing on inference time, Floating-Point Operations (FLOPs), and memory consumption. All models are tested on identical hardware setups: an NVIDIA RTX 3090 GPU and an Intel Xeon CPU, with batch size fixed to 1 to simulate single-query inference typical of interactive systems. Table 7 summarizes the key efficiency metrics averaged over 1000 inference runs per dataset. From Table 7, HonestVQA demonstrates an average inference latency of approximately 119.6 ms per query, which is about 6%-20% slower than the baseline models due to the additional uncertainty calibration and contrastive enforcement modules integrated into the architecture. Despite this, the latency remains within acceptable limits for real-time applications, especially in contexts where reliability and ethical correctness outweigh marginal speed gains. Regarding computational complexity, HonestVQA requires on average 69.4 GFLOPs, representing a 7%-18% increase relative to baselines. This overhead is primarily attributable to the dual loss functions enforcing confidence-accuracy alignment and ethical contrastive constraints, which necessitate additional forward and backward computations during inference calibration. Memory usage also increases modestly to 3075 MB on average, about 5%-14% more than competing models. The memory footprint remains manageable on modern GPUs, supporting the feasibility of deploying HonestVQA on standard deep learning infrastructure without requiring specialized hardware.
@@ 7 Conclusion and Future Works
In this work, we introduced HonestVQA, a novel framework that integrates uncertainty-aware alignment and contrastive ethical enforcement to effectively detect hallucinations in DocVQA systems. Through comprehensive experiments on diverse datasets—we demonstrated significant improvements in answer correctness, calibration, and cross-domain generalization compared to strong baselines. Our ablation studies confirmed the complementary role of each component, and efficiency analyses showed HonestVQA practical feasibility. Future research will focus on advancing domain adaptation techniques to further enhance robustness across unseen data distributions, and exploring lightweight calibration modules for deployment on edge devices. Additionally, incorporating user feedback for interactive hallucination correction and extending the framework to multimodal dialogue systems represent promising directions to improve DocVQA reliability and ethical safety.
@@
This work aims to enhance the trustworthiness and ethical reliability of DocVQA systems by reducing hallucinated and potentially misleading answers. HonestVQA promotes transparency through uncertainty-aware calibration, encouraging responsible AI deployment. We acknowledge the risk that no model can be entirely free of errors or biases, especially when applied across diverse real-world scenarios. Thus, we emphasize that HonestVQA is intended as a tool to assist, not replace, human judgment, particularly in high-stakes contexts. All datasets used comply with their respective licenses, and no private or sensitive data was involved. We encourage further research on fairness, bias mitigation, and inclusivity to ensure equitable AI systems, and advocate for ongoing monitoring of model outputs to safeguard against misuse or unintended harm.

### v2

@@ 5.2 Cross-Domain Generalization Testing
To assess the robustness and generalization ability of HonestVQA framework, we conduct a series of cross-domain testing experiments. These experiments evaluate whether the hallucination detection model trained on one dataset can effectively identify hallucinations in DocVQA outputs on different datasets. Such generalization is crucial in real-world scenarios where the distribution of questions and visual-textual content varies significantly across domains such as scanned documents, infographics, and structurally diverse textual scenes. We train HonestVQA on the source dataset and evaluate it on a different target dataset without any further fine-tuning, measuring hallucination detection performance using Accuracy and F1-score as the primary metrics. From Table 3, we observe that HonestVQA generalizes robustly across domain shifts. Notably, when trained on InfographicsVQA and evaluated on SpDocVQA, the model achieves a high F1-score of 76.1%, outperforming the reverse setting (SpDocVQA $\rightarrow$ InfographicsVQA), which yields 71.8%. This suggests that the high-density, information-rich visual patterns in InfographicsVQA provide transferable inductive biases that enhance hallucination detection in more structured domains like documents. Similarly, the SROIE $\rightarrow$ SpDocVQA setup results in an F1 of 72.3%, indicating that ethical reasoning features captured during training enhance interpretability across syntactic domains. However, the model exhibits relatively lower performance when transferring from SpDocVQA $\rightarrow$ SROIE (F1 = 69.8%), highlighting the challenges posed by ethical hallucination detection under unfamiliar structural constraints. Nonetheless, the use of uncertainty-aware calibration via confidence-alignment and contrastive ethical loss contributes to soft regularization of decision boundaries, allowing for improved generalization even in low-overlap semantic settings. Interestingly, models trained on SROIE also generalize well to visually and semantically distinct domains such as InfographicsVQA, achieving 67.2% F1. This supports our hypothesis that the inclusion of contrastive ethical supervision enforces more generalizable representations. Furthermore, the relative drop in performance in domain transfer settings (typically within 4%–6% of in-domain results) underscores the importance of calibration-aware models in mitigating performance degradation due to domain shift.
@@
Table 5: Sensitivity analysis of alignment and contrastive hyperparameters on SpDocVQA, InfographicsVQA, and SROIE datasets. Note: Accuracy (Acc) and H-Score (H) are used as abbreviated.
@@ 5.3 Ablation Study
To thoroughly evaluate the contribution of individual modules in HonestVQA, we conduct an ablation study across three datasets by systematically disabling the confidence-accuracy alignment loss and the contrastive ethical enforcement loss. Table 4 shows that the removal of either module consistently degrades performance across all datasets in terms of accuracy and H-Score, confirming their complementary roles. For instance, on SpDocVQA, the full model achieves 75.9% accuracy and an H-Score of 0.113. Removing the alignment loss drops accuracy to 72.1% and worsens H-Score to 0.172. On InfographicsVQA, the full model yields 68.3% accuracy and an H-Score of 0.134, whereas removing contrastive enforcement lowers accuracy to 65.2% and degrades H-Score to 0.189. Similar trends are observed on SROIE, where the full model achieves 80.2% accuracy and 0.096 H-Score, significantly outperforming the ablated variants.
We further conduct a hyperparameter sensitivity analysis on alignment weight $\alpha$, contrastive margin $m$, and loss weights $\lambda_{1}$ and $\lambda_{2}$ across three datasets and summarize the trends in Table 5. The model maintains high performance for $\alpha$ between 0.5 and 1.5, margin $m$ from 0.3 to 0.7, and loss weights $\lambda_{1}=0.1$, $\lambda_{2}=0.05$. Deviations outside these ranges result in decreased accuracy or calibration degradation.
@@ 6.2 Computational Analysis
We evaluate the efficiency of HonestVQA against baseline DocVQA models in terms of inference latency, FLOPs, and memory usage. All models are tested using an NVIDIA RTX 3090 GPU and Intel Xeon CPU with a batch size of 1 to simulate real-time settings. As shown in Table 7, HonestVQA incurs an average latency of 119.6 ms per query—6%–20% slower than baselines—due to uncertainty calibration and contrastive modules. It consumes 69.4 GFLOPs (7%–18% higher) and 3075 MB memory (5%–14% higher). Despite the overhead, it remains deployable in real-time systems where ethical reliability is critical.
(a) Ethical Risk Heatmap Confidence vs Uncertainty
(b) Semantic Drift under Ambiguity
(c) Contrastive Embedding Separation
(d) Attention IoU over Epochs
Figure 1: HonestVQA enhances ethical calibration, semantic stability, embedding separation, and multimodal grounding through uncertainty-aware learning,
6.3 Qualitative Analysis
HonestVQA improves ethical alignment by reducing overconfidence in uncertain scenarios, as seen in the risk heatmap (Fig. 1a). Semantic drift under ambiguity is mitigated, with more stable embeddings (Fig. 1b). Contrastive embedding separation (Fig. 1c) shows clearer distinction between aligned and misaligned responses, supporting improved representation learning. Finally, Fig. 1d shows consistent attention patterns over epochs, highlighting better multimodal grounding and interpretability.
@@ 7 Conclusion and Future Works
In this work, we introduced HonestVQA, a novel framework that integrates uncertainty-aware alignment and contrastive ethical enforcement to effectively detect hallucinations in DocVQA models. Through comprehensive experiments on diverse datasets—we demonstrated significant improvements in answer correctness, calibration, and cross-domain generalization compared to strong baselines. Our ablation studies confirmed the complementary role of each component, and efficiency analyses showed HonestVQA practical feasibility. Future research will focus on advancing domain adaptation models to further enhance robustness across unseen data distributions, and exploring lightweight calibration modules for deployment on edge devices. Additionally, incorporating user feedback for interactive hallucination correction and extending the framework to multimodal dialogue systems represent promising directions to improve DocVQA reliability and ethical safety.
@@
This work aims to enhance the trustworthiness and ethical reliability of DocVQA models by reducing hallucinated and potentially misleading answers. HonestVQA promotes transparency through uncertainty-aware calibration, encouraging responsible AI deployment. We acknowledge the risk that no model can be entirely free of errors or biases, especially when applied across diverse real-world scenarios. Thus, we emphasize that HonestVQA is intended as a tool to assist, not replace, human judgment, particularly in high-stakes contexts. All datasets used comply with their respective licenses, and no private or sensitive data was involved. We encourage further research on fairness, bias mitigation, and inclusivity to ensure equitable AI models, and advocate for ongoing monitoring of model outputs to safeguard against misuse or unintended harm.
Acknowledgments
This research was supported by Jamia Hamdard University, New Delhi, India and funded through the Visvesvaraya Fellowship, Government of India. Furthermore, this research is partially sponsored by funding from Yonghua Foundation.